OpenStudy (anonymous):

how can x^2>x^4>x^6.......when x belongs to (-1,1) ?

OpenStudy (anonymous):

I am not sure exactly what you are asking, however when x = -1, x^2 = 1, x^4 = 1 and x^6 = 1... when x = 0 each of those would = 0 and when x = 1, each of those would equal 0.. HOWEVER, in between those values... like -1/2 or 1/2 x^2 = 1/4 which is greater than x^4 = 1/16 which is greater than x^6 = 1/64. This would be true for all fractions between -1 and 0 and 0 and 1...

OpenStudy (anonymous):

it's not true for -1 and 1= both will have same values, so take numbers in between. like +- 1/2 or +- 1/4 and you will see the difference. :)